### v1

# Kaiwu: A Multimodal Manipulation Dataset and Framework for Robot Learning and Human-Robot Interaction

Abstract
Cutting-edge robot learning techniques including foundation models and imitation learning from humans all pose huge demands on large-scale and high-quality datasets which constitute one of the bottleneck in the general intelligent robot fields. This paper presents the Kaiwu multimodal dataset to address the missing real-world synchronized multimodal data problems in the sophisticated assembling scenario, especially with dynamics information and its fine-grained labelling. The dataset first provides an integration of human, environment and robot data collection framework with 20 subjects and 30 interaction objects resulting in totally 11,664 instances of integrated actions. For each of the demonstration, hand motions, operation pressures, sounds of the assembling process, multi-view videos, high-precision motion capture information, eye gaze with first-person videos, electromyography signals are all recorded. Fine-grained multi-level annotation based on absolute timestamp, and semantic segmentation labelling are performed. Kaiwu dataset aims to facilitate robot learning, dexterous manipulation, human intention investigation and human-robot collaboration research.
Index Terms:
Robot learning, Embodied AI, Human-Robot Interaction, Multimodal Fusion
I Introduction
Robots represent a key platform for artificial intelligence (AI) and are often considered the primary carrier of Embodied AI [1]. With the shift in demand from cyberspace to physical environments, traditional industrial robots are inadequate to adapt to changing environments. There is a growing need for intelligent robots capable of interacting with humans in uncluttered environment [2] and master human-level skills. However, the path to attaining human-level skills remains challenging and is still an ongoing research problem.
Datasets are the fundamental part of robots and embodied AI, and cutting-edge methods, including imitation learning or foundation models, all pose significant and growing demands on the high-quality, large-scale and multi-modal datasets [3]. Imitation learning is primarily achieved through human demonstration data, and most of the methods are based on vision data [4] or combined with depth data. New challenges on the data arise from changes in target skills, shifting from simple, fixed motions required for specific tasks, such as picking up an object, to more complex motions. In addition, demonstration via teleoperation represents another alternative approach. For example, Finn et al. proposed mobile ALOHA [5] which used whole-body teleoperation data collection methods. Tesla proposed to use real-human motion capture data to train the Optimus humanoid robots [6]. To be noteworthy, with the emergence of foundation models [7] including ChatGPT, the importance of data volume and quality in determining the algorithm’s upper limits has become more widely recognized and there are a few pioneer works exploring the real-world datasets towards the foundation models for robotics.
RT Series [8], [9], [10] focus on long-term task samples that include multiple activity instances, emphasizing autonomous environmental perception and logical reasoning ability. The emerging challenges for foundational models now entail a shift from managing relatively short-term tasks to attaining long-term autonomy. Philosophically, long-term autonomy requires interaction with the environment and its context, because the brain-inspired algorithms alone are insufficient. Consequently, embodied AI addresses the interplay between humans, robots, and their environment.
Nevertheless, current datasets all suffer from the following critical limitations. First, most of the data mainly relies on the videos and thus lacks dynamics information. Computer vision methods in the form of images or videos, although advancing and widely deployed, can only represent kinematics information including trajectory and velocity. However, robot learning is a process of dynamics with the surroundings. Missing dynamics information including force will deteriorates the learning performance, resulting in the superficial learning. Second, there is a lack of a universal, sophisticated, intuitive human-level perception framework. The working environment for intelligent robots changes from fixed and clustered settings including manufacturing scenarios to open and complex settings including in situ homes. The environment not only involves static and dynamic objects but also includes human beings. Previous datasets have only incorporated certain sensing modalities, such as video, Inertial Measurement Unit (IMU), etc. which are inadequate to address the demands of increasingly complex tasks. Furthermore, one reason that the manipulation skills of robots cannot surpass those of humans is the lack of fundamental understanding of neural mechanisms. Additionally, relying solely on video as a modality is insufficient to fully reveal how humans accomplish such complex tasks.
To address the above challenges, this paper presents a multimodal framework and dataset for robot learning and evaluation. The key differences from previous datasets include the integration of multimodal data encompassing humans, robots, and the environment. This dataset not only supports robot learning from human demonstrations but also facilitates the prediction of human intentions to enable more effective policy development. We aim to develop skills sets for robot learning and the name Kaiwu is inspired by the ancient Chinese literature Tiangong Kaiwu (also known as The Exploitation of the Works of Nature), a comprehensive encyclopedia encompassing a wide range of fields, including agriculture and craftsmanship. The contribution of this paper is as follows:
(1) A multimodal data collection framework is proposed, featuring full situational awareness including manipulation dynamics information, human mannipulation neural signals and attention information and multi-view manipulation vision information. This framework aims at complex scenarios including human assembly, towards the universal manipulation ability for embodied AI.
(2) High-quality, large-scale multimodal data for long-horizon autonomy are collected, utilizing state-of-the-art ground truth techniques for fine-grained and complex manipulation process, holding potential for future benchmark.
(3) The dataset is accompanied by an annotation of spatio-temporal relationships. Rich and fine-grained cross-modal synchronization data annotation are performed, including 298 annotations of regions of interest as personal attention data, 536,467 annotations of closed area elements for image segmentation, 7,197 motion segmentation events for left and right hand dexterity manipulation, 4,467 gesture event annotations, and 4,959 annotations for gesture classification, which significantly enhanced the cross-modal learning and multimodal fusion capabilities and its interpretability.
The rest of the paper is summarized as follows. In Section 2-4, research on related datasets, the experimental setup, process design, and the formulation of the target problems are first introduced. In Section 5, a detail introduction to the annotation of Kaiwu dataset is presented. Then, the descriptive statistics and directory structure of the dataset are presented, discussed, and introduced by category in Section 6. API, project official website and repository, known issues, proposed improvements, and future research directions are discussed in 7-8 sections. Finally, the conclusion is drawn in Section 9.
Figure 1: A framework of wearable and environment-mounted sensors records rich activity information in the assembly environment.
II Related work
The main focus of this paper is on how to build a multimodal dataset for robot learning via human demonstration for human-level skills and future human-robot collaboration, and thus the related key technology are first reviewed.
II-A Robot learning dataset
Multi-modal input can serve as a crucial foundation for robots to learn human behavior, enhancing the accuracy and efficiency of the learning process. The dataset focuses on teaching robots to perform daily interactions in changing environments through human demonstrations [7]. JD ManiData, ManiWAV [11] and RH20T [12] are designed to facilitate the acquisition of diverse and generalizable skills by robots, and the robot hardware platform for DROID data [13] is collected through a distributed collection approach to improve the performance and generalization of the robot’s operational strategies. A large-scale benchmark dataset called GraspNet-1Billion is proposed by Fang et al.[14], aiming to improve robust object grasping in cluttered scenes under multimodal data conditions. The Open-X implementation [15] enables the integration of large-scale data volumes and validates the positive migration of data between robots to improve multi-robot capabilities. However the above datasets suffer from data homogenization. The ability to efficiently and purposefully train embodied intelligence models in complex scenarios is limited. Building on the excellent datasets above, more research has been performed to further extend on a particular aspect. The evolution of the RT series datasets exemplifies this progression. The Robotics Transformer 1 (RT-1) [8] is designed to achieve multi-task robot control in real-world environments by utilizing large-scale datasets and high-capacity model architectures. The RT-1 model effectively generalizes by incorporating datasets that encompass a wide range of robotic operations and environmental scenarios. The RT-1 model dataset comprises text instructions, robotic vision images, image frame actions, and commands. However, there are limitations in the selection of scenes and target objects. In the subsequent RT Series models, the datasets are continuously expanded in actions, target objects and target position aspects. RT-2 [9], building on the foundation of RT-1, utilizes Web VQA (Web-based Visual Question Answering) to match the image frames from RT-1 with corresponding text tokens, effectively addressing the limitations related to target objects and positions. The Open X-Embodiment dataset [15] proposes the integration of the RT-2 database with additional datasets containing various actions and skills. The initiative aims to train the RT-X model, which is designed to excel in versatility across diverse usage scenarios. In the latest RT-H, the dataset has been augmented with human-robot interaction data pertaining to human interventions, thereby enhancing the model’s learning capabilities. In addition, the ARIO [16] dataset is optimized based on the Open X-Embodiment dataset. It contains more sensory data in 5 modalities: 2D, 3D, text, tactile, and sound, and covers two types of tasks: operation and navigation. It contains both simulation data and real scene data, and includes a variety of robot hardware subjects, retaining a large amount of data and a uniform format. However, These datasets are limited by insufficient dynamic kinematic information and a lack of authenticity in the raw data, and suffer from the issue of labeling generality.
II-B Human activity recognition dataset
Enabling robots to accurately recognize and understand human activity is an essential aspect of developing embodied intelligence. Therefore, high-quality datasets are urgently needed. The industry-oriented dataset [17] is a dataset of human movement in industrial activities and is intended to be used to classify, predict or evaluate human movement in an industrial environment. The ActionSense dataset proposed by DelPreto et al. [18], which focuses on kitchen scenes, provides a multimodal human activity dataset. Yoon et al. [19] proposed HUMBI which is a multiview camera dataset for human body modeling. The dataset addresses the limitations of human expression datasets by modeling human appearance in high resolution. The HUMBI dataset is not only significant for appearance modeling but also serves as a powerful multi-view benchmark for establishing more robust and generalizable human body reconstruction and recognition models. This, in turn, enhances the identification capabilities of agents. The Toyota Smarthome Untrimmed (TSU) dataset proposed by Dai et al. [20] collects untrimmed video data of spontaneous activities in home settings. This untrimmed data can help robots understand the causality of human activities before and after, thereby enhancing their ability to recognize human behaviors. The dataset [21] is designed to help autonomous robots become aware of the overall intent of the team they are working with in order to effectively help their human counterparts or augment the team’s capabilities. The current research lacks an extension of the application scenarios for the dataset and a focus on assembly task related scenarios. In addition, there is a lack of balance on data modality homogeneity, cross-modality temporal consistency, and task causality analysis.
II-C Human robot collaboration dataset
Embodied intelligence emphasizes the interaction capabilities of intelligent systems with their physical bodies and environments. Human-robot collaboration enables agents to learn how to better adapt to and comprehend their surroundings through interactions with humans, thereby enhancing the generalizability and stability of behavioral patterns. Multimodal datasets integrate various data types, such as text, images, sound, and video, which contain rich information about user intentions and emotions. In human-robot interaction, multimodal data can assist agents in better understanding the needs of operators, improving accessibility and information perception capabilities, and providing a more natural and intelligent mode of interaction. The dataset [22] focuses on teaching robots to perform daily interactions in changing environments through human demonstrations. The RBO dataset [23] records data on human interaction with articulated objects, which provides a valuable data resource for robotics and inspires the construction of subsequent human-robot interaction datasets. The RT-H [10] method adds interventions of human-robot interaction behaviors to the RT-2, which enhance the robot’s understanding of interaction action or language by describing motions with more fine-grained phrases. Thereby, RT-H builds an action hierarchy, enabling the agent to learn from the interventions of human language motions.
Additionally, eye gaze, EMG signals, IMU and skeletal postures contain a wealth of information about human operational intent that can be utilized to enhance the accuracy and diversity of collaborative human-robot interactions. Mapping the region of interest using eye-tracking devices along the viewing trajectory can help the robot understand the human’s intentions during the assembly process. EMG and IMU signals record the behavior patterns of human arm muscles during the execution of an action. This data can help the robot learn and predict hand gestures to respond to human behavior. The HARMONIC dataset [24] constitutes a comprehensive collection of human-robot interactions within a shared autonomous setting. Throughout the data acquisition process, the robot executes a series of experimental maneuvers facilitated by a joystick which is controlled by a person with wearable sensors. This approach enables the aggregation of data that is rich in human intent, thereby providing valuable insights into the nuances of human-robot interaction. Consequently, the HBOD dataset [kang2023hbod] employs more wearable motion sensors to capture intricate details of human movement and direct interactions with an array of tools. This methodology significantly enhances the robot’s capacity to comprehend human movement intentions and strategic maneuvers within a specified context. The dataset [25] facilitates an expansive discourse on overarching learning challenges within the realm of hand-object interaction, encompassing aspects such as intention recognition and motion generation, with implications for the field of Robotics. The OAKINK2 dataset [26] provides multi-view image streams and accurate pose annotations of the human body, hands, and various interacting objects. This rich data supports a variety of applications such as interaction reconstruction and motion synthesis.
The aforementioned datasets employ indirect methods of data collection, which leads to deficiencies in recording and describing dynamic information. The action interactions within the designed scenarios fail to connect motion units into a narrative process, lacking in causality and coherence. In Kaiwu dataset, wearable sensors are utilized to collect dynamic and static data directly. The scenario design integrates assembly action units into the entire assembly process, rendering the human-robot interaction more narrative and causally connected. Furthermore, the interaction process is complemented with the establishment of a coordinate framework for experimental environment, thereby enriching the dynamic information and spatial integrity of the interaction. A detailed comparison of the proposed methods with other SOTA datasets was performed (see Table I).
TABLE I: A Comparison of related datasets. Motion capture contains 3D skeleton and ground truth.
Dataset Modalities Environment/Activities
TSU RGB,Depth,3D Skeleton Daily actions
Harmonic Gaze,EMG,RGB,Depth Meal
Hbod 3D Skeleton,Tactile,Hand Pose,IMUs Tool Operation
Humbi RGB,Depth, 3D Skeleton Body Expression
OXE Mainly RGB,Depth Multiple Scenarios
Actionsense IMUs,3D Skeleton,Hand Pose,Gaze Kitchen Activities
EMG,Tactile,RGB,Depth,Audio
Kaiwu IMUs,Motion Capture,Hand Pose,Arm,Gaze Industrial Assembly
EMG,Tactile,RGB,Depth,Audio
III Data collection platform setups
The data collection platform facilitates the synchronized streaming, storage, and visualization of info from wearable and environment-mounted sensors with varying sampling rates and data formats. The platform’s structure simplifies the integration of newly added sensors. The multi-threading and multi-processing is designed to ensure efficient CPU and RAM usage during the collection. Additionally, it offers post-processing capabilities that allow for the combination of streamed data with data processing by third-party applications.
The platform also aims to facilitate efficiency of data collection with fewer experimental artifacts by allowing researchers to operate, record metadata without tedious post-processing, and pay more attention to the process of subject build.
TABLE II: Overview of sensors setups.
Device Sensor Type Data Sreams Sampling rate [Hz] Calibration Third-Party Recording Software
Tactile sensors grip force feedback
WISEGLOVE19FE Angle sensors finger angel 100 hand pose calibration GraspMF
Arm IMU hands,arm quaternion
Trigno Biofeedback System EMG , IMU EMG signal,ACC 4,000 Stand with known locations and poses —————————————
First-person camera First-person videos Tobii Lab
Tobii Pro Glasses 3 Infrared eye camera gaze point 50/100 Pupil Capture calibration Tobii Pro Eye Tracker Manager
IMU pupil details
Nokov XINGYING Motion capture camera 340 Stand with known poses and locations XINGYING system 2.1.0
Color videos
Azure Kinect DK RGB+D raw-format frame images 60 Place with participant in field of view —————————————
Depth data
Microphone Omnidirectional,cardioid Raw audio recordings 48,000 Place in preset position —————————————
III-A Application and use cases
The main purpose of this dataset is to provide rich information on how human can achieve dexterous operation and how these information can help robot acquire human-level intelligence. Therefore, we identify several critical processes of human dexterity including dynamics information during the manipulation, attention during the cognition, understanding of muscle group mechanism and assembly logic, ubiquitous environment information along with the operation process and also ground truth recording. The overview diagram is shown in Fig. 1.
III-B Sensor setups
Targeted at the above requirements, cutting-edge sensing acquisition equipment are utilized. These devices include wearable sensors and environment-mounted sensors (Fig. 1).
III-B1 Data glove
TABLE III: Data format of glove data.
sequence column 1 $sequence$
timestamp column 2 $timestamp$
hand column 3-6 $[hand_{w},hand_{x},hand_{y},hand_{z}]$
forearm column 7-10 $[fore_{w},fore_{x},fore_{y},fore_{z}]$
upper arm column 11-14 $[up_{w},up_{x},up_{y},up_{z}]$
angle sensor column 15-33 $[thumb,thindex,index,inmid,middle,$
$midring,ring,ringlittle,little,thumbcmc]$
tactile sensor column 34-52 $[p0-p18]$
Data gloves are utilized to simultaneous capture the hand movement and palm tactile interaction information. Therefore, the WISEGLOVE glove (Fig. 2) is chosen which can drive 3D animated human hand movements in real time. 19 finger angle sensors and 19 finger pressure sensors are equipped with a sensing accuracy of 9g, which fulfills the spatial and accuracy resolution of the experimental requirements. Additionally, there is an arm sensor that recorded quaternion data from the palm, forearm, and upper arm sensors. The system can capture human hand motion in real-time and drives 3D animation simultaneously to depict hand motions. This system is used to study participants’ hand movements and finger coordination during assembling tasks.
Figure 2: Data glove with angle sensors and force sensors.
III-B2 EMG and ACC
TABLE IV: Data format of EMG (or ACC) data.
index sequence row 1 data index
timestamp row 2 timestamp of data
left hand row 3-10 raw data of EMG or ACC sensors
right hand row 11-17 raw data of EMG or ACC sensors
The muscle activity of the participants is accurately collected for studying the relevance of EMG data during assembly tasks. The Trigno Sensors are equipped with an integrated 9-degree-of-freedom IMU capable of transmitting data pertaining to acceleration (ACC). This information can be used to discern movement activity that is time-synchronized with the EMG signals. They monitor the electrical signals of the muscles through electrodes attached to the participant’s skin and converts them into readable digital data. A total of 16 EMG sensors are used, with 8 sensors attached to the left side and 8 to the right side, each being attached to the basic muscle groups of the forearm.
III-B3 Environment depth and visual information
Kinect is positioned directly in front of the participant to record the environmental information including RGB and depth information (Fig. 3) during the assembly process. This camera is used to capture participants’ body postures and movements in real-time for investigating their motor performance and movement control in specific tasks.
Figure 3: RGB-D information.
III-B4 Eye tracker
Eye-tracking data, including participants’ gaze points mobility, are obtained to invesitgate visual attention and cognitive processes. The Tobii Eye Tracker (Fig. 4) is a binocular stereoscopic dark pupil tracking device that relies on corneal reflexes to record and analyze human eye tracking and first-person information in assembly processes.With an average accuracy of 0.6°, an average precision of 0.03°and an average data loss rate of 0.01%, the Tobii Pro Glasses 3 guarantees quality data collection.
Figure 4: Attention information of eye tracker.
III-B5 Environment sound information
The microphones are used to capture ambient sound as a complement to modality, such as the sound of a tool being placed and rubbing against parts. In the experiment, four microphones are placed in different locations. Numbered 1-2, are placed on the operator’s table to record the sound information of the operating environment. Microphones numbered 3-4 are placed in the accessory area to record ambient sound information for gripping tools and parts.
III-B6 Ground truth
The Nokov Optical 3D Motion Capture System is used to capture and record the movement and actions of a person or object in three-dimensional space. It utilizes high-precision optical sensors and cameras to track the movement trajectory of the object being measured and then transmitting the data to a computer for analysis and rendering. In the experiment, participants are fitted with reflective markers at 37 key points on the body to calibrate and build a mannequin. With this system, the ground truth information for analysis can be accurately captured. The location of reflective feature points, camera placement, and human skeleton modeling of the motion capture system are shown in Fig. 5.
Figure 5: Motion capture system calibration.
Figure 6: Overview of assembling process. C8 and C14 have different focuses. The former collected experimental data on columnar parts that participants tightened by hand in a situation where there is ample mounting space. The latter captures experimental data by tools in a confined space with limited hand movement.
III-C Calibration
Initial calibration of different participants can significantly improve the quality of metadata.
III-C1 Equipment initialization calibration
The human eye exhibits natural individual differences in shape and geometry. The location of the visual central fovea, responsible for perceiving the details and colors of a focused target, varies from person to person. To address individual differences in eye tracking performance, the Tobii Pro eye tracker utilizes a one-point calibration procedure to enhance the eye tracker’s accuracy in predicting the subject’s visual position. To ensure that accurate eye-tracking data is obtained in an experiment, the eye-tracker must be calibrated for each subject. The data recorder positions the calibration card facing the subject, ensuring that the card is held at an arm’s length away from the participant. The participant should wear the eye-tracker’s head-mounted module and maintain focus on the center of the calibration card throughout the calibration process.
Physical stature varies across participants, so for each collection experiment, the motion capture system needs to be calibrated to the skeleton of the participants after the reflective markers are installed. The participant’s arms are opened at shoulder level and slowly rotated to ensure that all reflective markers on the participant’s body are recorded in the system (Fig. 5).
When performing arm activity manipulations, the active muscle groups in the lower arm vary from participant to participant. Therefore, it is necessary for the participant to continually rotate the wrist and measure the most significantly active muscle groups to place the EMG acquisition unit to confirm that the most distinctively characterized EMG signal data is captured.
III-C2 Process initialization calibration
After starting the experiments, the participants perform calibration of gesture movements to activate data synchronization and initialize the 19 sensors on the gloves. Each group of experiments can only be performed after completing the calibration process mentioned above for the corresponding assembly operation to ensure the synchronization of data acquisition.
III-C3 Data synchronization
The data collection platform with modular programming and ultimately implemented multi-threading is designed to ensure simultaneous start and synchronized acquisition of each acquisition module. The platform ensures the independence of data collection and the convenience of code adjustment. For different devices with different sampling rate in Subsection 6.1. Absolute timestamp records are used during data collection, and the output of each module’s data stream is synchronized through absolute timestamps.
IV Data collection protocol
IV-A Participants
Twenty participants with a mean height of 172.4$\pm$6.5 cm and a mean age of 23.95$\pm$5.1 yrs are recruited in this process, and all the consents are given. Each volunteer is recorded for 2 hours. The entire data collection process span 11 working days. Approval of all ethical and experimental procedures and protocols is granted by Institutional Review Board of Tongji University under Application No.TJDXSR055, 2023. The experiments are performed in the Kaiwu Lab of Tongji University.
IV-B Assembling process
After the participant completed donning the device, calibration session, and data stream synchronization, data acquisition of the assembly actions began sequentially. During a single action acquisition session, the process would last 140 seconds, involving 0-10 seconds of data synchronization to receive calibration and 130-140 seconds of buffer time.
IV-B1 Assembly activity
In the assembly process of the robotic arm, links that closely reflect human dexterity and dynamics are selected for data collection. The entire assembly process is divided into 15 assembly links, with each link divided into left and right fine-grain motion elements labeled as action units, which are recorded in glove exported model playback. The segmentation facilitates the coding and integration of the entire assembly process into an action ensemble, and also facilitates subsequent labeling and recall. The assembly activities are shown in Fig. 6.
IV-B2 Assembly equipment and parts
For the above action set, the relevant set of frontal assembly objects includes the robot front arm, robot end arm, robot middle arm, robot arm base, rear arm part, flange bearings, linkage, motor, steering bearings, brass post, steel post, bolt spacer, bolts, nuts, and power cable.
Instead of subdividing the action units into finer degrees in our experiments, we remain at the level of a relatively large patchwork of components. This is due to the fact that it is difficult for the finger tactile sensors to detect changes of data in small and delicate components. Before proceeding with data collection, an initial pre-assembly of the original small components is conducted to meet the acquisition requirements.
IV-B3 Tools
For all divided action sets, the following execution tools are required: drill, snips, file, polisher, vernier calipers, sharp-nosed pliers, and screwdriver.
V Data annotation
An initial labeling of the dataset is performed to facilitate potential robot learning and human-robot interaction tasks. Various data are labeled to enhance the usability of the dataset. Subsequent users have the flexibility to annotate data according to their specific project requirements, while our initial annotations serve as a fundamental reference paradigm (Fig. 7).
Figure 7: Overview of annotation.
The annotation information, such as labels and instance, is shown in Table V.
TABLE V: Overview of annotation.
type of mission annotation element object tags instance
gesture classification picture 10 4959
AOIs video 30 298
semantic segmentation closed area 30 610778
action segmentation video clip 26 7197
gesture segmentation video clip 9 4467
V-A Action&gesture segmentation and export
Figure 8: Label distribution.
Fine-grained segmentation is conducted based on the timestamp series of actions and gestures to identify the start and end timestamps of the multimodal data collected from each assembly task action unit. The annotation includes:
  • Action-level Annotation: Coarse segmentation of left and right hand actions.
  • Fine-grain Gesture Annotation: Within the coarsely segmented time interval, a fine-grain segmentation of the left or right hand states is made
For action segmentation, the left and right-handed action segments of each participant are divided as a whole into time intervals for the stage actions. And the free gaps between movements are retained to preserve the causality and coherence between movements. The segmentation of movements includes basic and specific movements, with the former including generalized movements such as approaching, grasping, placing, and tightening screws. The latter includes actions specific to each assembly-specific task, such as the wiring action in the motor mounting session.
The execution of actions in each segmentation interval is transitioned by the change of several gesture states, the actions of the left and right hands are further divided into corresponding hand state intervals, which are notated as Fine-grain Gesture Annotation. In the division of hand states, all hand states are divided into 8 gesture states [27]: Cylindrical grasp (Cyl), Oblique palmar grasp (Obl), Lumbrical grasp (Lum), Intermediate power-precision grasp (IntPP), Pinch grasp (Pinch), Lateral Pinch (LatP), Special pinch (SpP), Non-prehensile grasp (NonP).
Firstly, according to whether or not the palm is involved, then they are divided into Group1: Cyl, Obl, IntPP and Group2: Pinch, LatP, SpP, Lum. in Group1 is further divided by the state of the thumb and the four fingers, which corresponds to Cyl: the thumb and the four fingers are naturally curved; Obl: the thumb is straight and the four fingers are curved; and the others are stipulated as IntPP. in Group2 is further divided by the number of fingers involved in the The number of fingers involved in the activity is further classified as Group2a: a small number (2-3) of fingers involved, containing Pinch and LatP; Group2b: four fingers involved, containing Lum and SpP.Finally, by the presence or absence of the lateral fingers, it is classified as LatP and SpP with lateral finger involvement, and Pinch and Lum without lateral finger involvement. The final label distribution is shown in Fig. 8.
In summary, for each time sequence, both action segmentation and gesture segmentation dimensions are included. For the labeled timestamped sequences, two annotation-level timestamped sequences in the glove data, EMG, and ACC signal data are derived to achieve absolute timestamped labeling of the EMG-ACC-Glove data series.
Figure 9: Semantic segmentation.
V-B Semantic segmentation
Semantic information annotation of 30 key objects in RGB video data is performed. By annotating the environment information from Kinect, which records the third-person view of the participants, the dataset is enriched with semantic information following the assembly logic. This enhancement aims to improve the robot’s logical understanding of the assembly process and human action learning. Labels and instances of annotation are shown in the Fig. 9.
In semantic segmentation, semantic information is labeled by pumping the captured RGB video at one-second intervals to obtain frames with corresponding timestamps. The participants and all interaction tools are labeled. For objects with multiple entities, each object is labeled followed by the corresponding index for differentiation. For objects that are partially blocked in use, the remaining portion is locally labeled. All labeling errors are controlled within 8 pixels.
V-C Gesture classification
According to the model playback, the state of the hand can be categorized [28] according to whether the thumb and four fingers are flexed or extended. Compared to the gesture segmentation in subsection 5.1, this part of the state segmentation simplifies the further division of the four-finger state in subsection 5.1.
V-D Area of interest (AOIs)
The exported first-person view videos are labeled with a region of interest based on key object manipulation. This region contains the focus and intentional expression information within the participant’s field of view, giving more weight to the data of objects in that specific area.
VI Data format
VI-A Descriptive statistics
A total of $6\times 15\times 20$ sets of experimental data on the assembling process are collected for Kaiwu dataset. The data comes from wearable sensors worn by 20 participants. Each participant is asked to complete 15 typical actions specified in the assembly process. Each participant spend approximately 19 minutes, resulting in the dataset representing about 6.3 hours of the assembling process. A summary of the related data is presented in Table VI.
TABLE VI: Descriptive Statistics.
Data Type Documentation Space Sampling Rate
Glove Data 264 MB 100 Hz
Glove Export 1,124 MB 20 Hz
Eye Tracking 14 GB 25 Hz
RGB-D Video 3,476 GB 60 Hz
Motion Capture Data 4,160 MB 60 Hz
Audio Data 7,955 MB 50 Hz
ACC Data 354 MB 40 Hz
EMG Data 362 MB 40 Hz
Figure 10: Kaiwu Dataset Structure Catalog
VI-B Detailed information
This dataset contains experimental data from 20 participants in 20 folders with subject number. The folder corresponding to each participant contains the collected data from 6 recording devices. The directory structure of the dataset is shown Fig. 10.
VI-B1 EMG data
The EMGData folder contains the ACC data output by the built-in IMU sensors and the EMG files measured and output by the EMG sensors. The ACC or EMG data are output by each of the built-in IMUs or EMG sensors respectively in 17 sets of data recorded as TIME.csv files. Each file contains 1 row for the calibration standard value and 2-17 rows for the ACC or EMG signals from the 16 sensors. The EMG_data directory structure is shown in Fig. 11.
{forest}
forest scheme [EMG or ACC data [ACC [C1-source_data [TIME.csv] ] [⋮] [C15-source_data] ] [EMG [C1-source_data [TIME.csv] ] [⋮] [C15-source_data] ] ]
Figure 11: Structure of the EMG or ACC data.
VI-B2 Glove data
For the assembly actions of C1-C15, the data gloves are transferred and stored as r/l_TIME.csv files with columns 1-12 representing the quaternion data of the palm, forearm, and upper arm, columns 13-31 representing the finger angle sensor data, and columns 32-50 representing the grip force sensor values. In addition, the data can be exported to present a visualization interface of the hand activity during data collection, and stored as an MP4 file. The GloveData directory structure is shown in Fig. 12.
{forest}
forest scheme [glove data [glove_data [left [l_TIME.csv (C1)] [⋮] [l_TIME.csv (C15)] ] [right [r_TIME.csv (C1)] [⋮] [r_TIME.csv (C15)] ] ] [glove_export [C1.mp4] [⋮] [C15.mp4] ] ]
Figure 12: Structure of the glove data.
VI-B3 RGB-D data
RGB-D data is stored in the C1-C15.mkv files in the RGB_video file and contains the timestamps used to generate the picture data, recorded as .txt files. The data generated based on the timestamp is stored in the IMG_picture file, including RGB images (.jpg file and .pcd file), depth information (.png file). The environmental information of the assembly data of C1-C15 is stored under 15 folders respectively. The kinect_data directory structure is shown in Fig. 13.
{forest}
forest scheme [RGB-D data [IMG_picture [C1 [color.jpg] [depth.png] [pcd] ] [⋮] [C15_source_data]] [RGB_video [C1.mkv] [C1.txt] [⋮] [C15.mkv] [C15.txt] ] ]
Figure 13: Structure of the RGB-D data.
VI-B4 Ground truth data
After the initial body joint point calibration, the NOKOV system captures and records the 3D spatial coordinates of each marker point as the assembly process proceeds, outputting the data as a (.cap .trb .xrb) File. Visualization is achieved by reimporting the data. Ground truth data directory is structured as Fig. 14.
{forest}
forest scheme [skeleton data [BodyNUM.mars] [Remotetake1.cap] [Remotetake1-Unnamed.trb] [Remotetake1-BodyNUM.trb] [Remotetake1-BodyNUM.xrb] [⋮] [Remotetake15.cap] [Remotetake15-Unnamed.trb] [Remotetake15-BodyNUM.trb] [Remotetake15-BodyNUM.xrb] ]
Figure 14: Structure of the skeletons data.
VI-B5 Eye tracker data
The Tobii-data folder contains the first-person perspective information (tobii-data) and the eye-tracking information (tobii-export). The former contains the acquisition data from C1 to C15 and generates the corresponding .gz data compression files (eventdata records the events, gazedata records the eye-tracking sensor data, and imudata records the IMU measurement data, which are combined and displayed in the tobi-export file) and the visualization .mp4 file. The latter contains C1-C15 .mp4 with gaze point movements, and .xlsx files listing the data tables measured by the built-in gyroscope, magnetometer, accelerometer and eye tracker. Each header in the files is labeled with the corresponding data meaning. The directory structure is shown in Fig. 15.
{forest}
forest scheme [eye
tracker
data [tobii_data [C1_source_data [eventdata.gz] [gazedata.gz] [imudata.gz] [recording.g3] [scenevideo.mp4] ] [⋮] [C15_source_data] ] [tobii_export [C1 [C1.mp4] [Data Export.xlsx] ] [⋮] [C15 [C15.mp4] [Data Export.xlsx] ] ] ]
Figure 15: Structure of the eye tracker data.
VI-B6 Voice data
The microphone outputs environment sound information as .wav audio files and .txt timestamp data. Microphones numbered 1-2 record the ambient sound in the assembly area and 3-4 record the ambient sound in the accessory area. The audio data directory structure is shown in Fig. 16.
{forest}
forest scheme [voice data [micophione1 [C1_TIME.wav] [C1_TIME.txt] [⋮] [C15_TIME.wav] [C15_TIME.txt] ] [⋮] [micophione4 [C1_TIME.wav] [C1_TIME.txt] [⋮] [C15_TIME.wav] [C15_TIME.txt] ] ]
Figure 16: Structure of the vioce data.
VII Accessing the data
The data has been uploaded to ScienceDB with DOI number https://doi.org/10.57760/sciencedb.14937. The homepage of this dataset can be searched using the keyword ”Kaiwu” on ScienceDB. The following files are provided for download: A compilation of all the data Kaiwu Data Annotation (31.84GB), which action segmentation (963MB), contains the timestamp exported delsys_data with ACC and EMG and the gloves_data with action labels. The AOIs (17.9GB) contains the annotation results and annotation process files which that can be opened using Tobii Lab software. The gesture segmentation (973MB) contains the delsys_data and gloves_data with fine grain gesture labels. The semantic segmentation (11.6GB) contains the part1-4.zip with RGB video annotation results, and finally gesture classification (417MB), contains the hand state classification data(classification result.zip).
In addition, the dataset contains some missing data and additions to the raw data (action segmentation supplement and raw data directories). The action segmentation supplement (193MB) adds four sets of experimental data that are missing from the action segmentation. The raw data (167GB) contains the raw voice data, motion capture data, RGB-D data, glove exported video and eye tracker exported video. The annotated data is saved and encapsulated as a .csv file. The timestamps can be used as indexes to call the corresponding data sources of different sensors at the same time, and contain the data annotation information of the corresponding timestamps.
VIII Known issues
Owing to limitation of computational load, certain data output streams may have periodic interruptions. The coverage of available multi-modal data is recorded in the descriptive statistics (Table VI), to assess the percentage of data stream dropouts.
Since the data gloves are not equipped with sensors at the fingertips, although participants are informed to minimize the use of bare fingertip areas not equipped with sensors, a small amount of this phenomenon still occurred in the experiments, resulting in incomplete data collection for a small portion of the glove sensing data.
IX Conclusion
This paper introduces a data collection framework and platform to build Kaiwu dataset, which enable to improve understanding of human assembly activity logic and robot learning. Multi-modal wearable sensors are equipped to capture various egocentric data which contains rich visual and dynamics info. This is also coupled with environmental and global ground truth info to build an absolute spatial coordinate frame. An initial labeling of the dataset is also annotated to facilitate the training of deep neural networks with multi-modal data.
Future work can utilize the Kaiwu dataset to develop knowledge pathways for embodied learning and analysis, exploring topics such as cross-modal prediction, assembly logic sequence prediction, assembly task planning, and robot self-assembly. In addition, we hope the collection platform of the Kaiwu dataset can be utilized as a medium for robot behavioral learning and human-robot skill transfer. In addition, more advanced wearable sensors can be added to explore new collection modes flexibly to enable the dataset for various application scenarios, serving as the benchmark for the database of robot foundation models, and thus pave the way for general embodied intelligence.